Parth (parthkohli):

Challenge: When @lgbasallote left OpenStudy, he left us with 44 theoretical dollars. Nine users fight over the money. Lord @shadowfiend decides that all the people will get a distinct amount of money based on a lottery and it must be a natural number. In how many ways can we distribute this money amongst these 9 users?

OpenStudy (anonymous):

a+b+c+d+e+f+g+h+i=44 where a,b,c,d,e,f,g,h,i are distinct whole numbers

OpenStudy (anonymous):

we need to find No. of solutions?

Parth (parthkohli):

Yesh.

OpenStudy (anonymous):

0

Parth (parthkohli):

Nailed it: I couldn't think of a better problem. -_-

Parth (parthkohli):

And why?

OpenStudy (anonymous):

Minimum sum of nine distinct whole number =45

Parth (parthkohli):

Wow, perfect.

Parth (parthkohli):

It was so easy that even my teacup said "zero, you dumbo!"

OpenStudy (agent0smith):

Easier than it seemed at first... 1+2+3+4...+9 = $45

OpenStudy (anonymous):

no i think if we are seeing whole no.s we should include a zero

OpenStudy (anonymous):

am i right or wrong someone plz tell.

OpenStudy (anonymous):

plz @ParthKohli replyy

OpenStudy (anonymous):

W H Y????????????

Parth (parthkohli):

Fixed, sorry

OpenStudy (anonymous):

i did'nt follow it plz explain the previous comment

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

Oh sorry I should have said natural numbers

OpenStudy (anonymous):

because all the people gets distinct amount of money

OpenStudy (anonymous):

ok i got it in the question there is natural numbers sorry ............at last

OpenStudy (anonymous):

Another question If 0 was included then how many ways?

Parth (parthkohli):

Then we actually have 8 people to give money to in natural numbers :-D 1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 = 36 So there are some ways.

OpenStudy (anonymous):

0+2+3+4+5+6+7+8+9=44 So, 9! ways

Parth (parthkohli):

Yeah.

OpenStudy (anonymous):

no in this case i think it can also be 0+1+2+3+4+5+6+7+16 we have to work out a hectic solution

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

but still a salute to your knowledge u people are too knowledgeous

OpenStudy (anonymous):

u r very correct

OpenStudy (anonymous):

i just mentioned 1 possibility it wud be kind enough if u two plz work out the solution and frankly speaking i don't so much calculations thanks

OpenStudy (anonymous):

0+2+3+4+5+6+7+8+9 0+1+3+4+5+6+7+8+10 0+1+2+4+5+6+7+8+11 0+1+2+3+5+6+7+8+12 0+1+2+3+4+6+7+8+13 0+1+2+3+4+5+7+8+14 0+1+2+3+4+5+6+8+15 0+1+2+3+4+5+6+7+16 So, 8*9!